# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ II.7 Spin-glass simulations and dedicated computers
Interest from computer scientists arises because finding the ground state of a spin glass is an archetypically hard computational problem. Indeed, if the interactions of the spin glass can only be drawn in a non-planar graph, [^1] A planar graph is one that can be drawn on a piece of paper without edge crossings, such as the square lattice with open boundary conditions. The difference between planar and two-dimensional is important, as discussed below. then finding the ground state belongs to the category of NP-complete problems [40, 155]. Should a truly efficient algorithm for finding the ground state of a non-planar spin glass be discovered, efficient algorithms for a large category of hard computational problems (such as the traveling salesman problem) would follow [236].
@@
Physicists are often more interested in the behavior of a system at finite temperature $T$ (one could say that physicists wish to minimize the free energy, rather than the energy). However, the physicist and the computer-scientist problems are tightly connected, because the free energy becomes the energy when we approach $T\!=\!0$.
@@
Difficulties faced by physicists depend strongly on whether they consider equilibrium physics or off-equilibrium dynamics:
@@
  • Equilibrium studies are (unfortunately) limited to small systems because reaching thermal equilibrium may require an inordinately long simulation time for some samples [in cubic lattices, equilibration times at low temperatures fluctuate from sample to sample by orders of magnitude, 12, 294, 109]. Further computational difficulties are posed by the fact that a large number of samples need to be simulated in order to get meaningful results, because quantities of interest wildly fluctuate from sample to sample.
@@
  • Non-equilibrium studies are conducted on lattices of large linear size $L$, sufficient to ensure that the coherence length $\xi(t,t_{\mathrm{w}};T)$ is always much smaller than $L$. This condition is required in order to compare with laboratory correlation lengths, which, in general, are smaller than physical sample sizes (with the exception of thin films, see Sec. V.1). Indeed, in space dimension $D$, the system can be approximated as composed of $[L/\xi(t)]^{D}$ more or less independent samples of effective size $L_{\text{eff}}\approx\xi(t)$. Hence, in close analogy to experiments, a simulation of just one very large sample automatically carries out the average over samples. The main difficulty is that one is limited to simulation algorithms that mimic natural dynamics, such as Metropolis [see, e.g., 262] and, as already explained in Sec. II.6, the natural growth of $\xi(t,t_{\mathrm{w}};H)$ is slow. This slowness calls for very long simulation times on few very large lattices.
@@
Simulating spin glasses is not always hard. Efficient algorithms for finding the ground state are known for planar interaction graphs. It is also important to note that planar and two-dimensional are not quite the same thing: the bilayer that was shown in Barahona [40] to pose an NP-hard problem is certainly two-dimensional but non-planar. The square lattice with periodic boundary conditions is mildly non-planar, but still allows one to find ground states for very large lattices [176, 76]. For $T>0$, the problem is also simpler on planar (or quasi-planar) graphs, such as the square lattice with periodic boundary conditions, for which non-local simulation algorithms (named cluster algorithms) can shorten simulation times by orders of magnitude [151, 286].
@@
Unfortunately, the situation is much worse for non-planar graphs, such as the simple-cubic lattice used in most $D=3$ simulations. At $T\!=\!0$, ground states have been found in cubic lattices with size $L\!\leq\!14$ with the help of a heuristic algorithm [199]. See Caracciolo et al. [76] for information on sure solvers. Many algorithms have been proposed for $T\!>\!0$: simulated tempering [198], umbrella sampling [see, e.g., 48, 156], the Wang-Landau algorithm [284] or population annealing [see, e.g., 196]. Nevertheless, the best-performing algorithm for $T\!>\!0$ equilibrium simulations appears to be parallel tempering [129, 152, 270], which, with the help of a custom-built computer, has been able to equilibrate cubic lattices of size $L\!=\!32$ at $T\!\approx\!0.64T_{\mathrm{g}}$ [13]. Temperature chaos, Sec. II.4, is the physical mechanism that makes it difficult for algorithms that exploit the correlation at different temperatures (e.g., simulated annealing or parallel tempering) to reach thermal equilibrium at even lower temperatures [111, 58]. [^2] In a very chaotic sample, thermalization achieved at a higher temperature would be almost irrelevant for sampling at a lower temperature. Thus, the thermalization process would need to restart at the lower temperature.
@@
Lacking better algorithms, the best we can do is to increase computing power. Computers specifically designed for high-energy physics simulations, such as RTN [22] and APE100 [42], have also been quite useful for spin glasses [see Badoni et al. [23], Ciria et al. [86] in the case of RTN and Marinari et al. [206, 205, 201] for APE100]. More recently, graphics processing units (GPUs) have proven to be extremely powerful for simulating quantum spin glasses [50]. Furthermore, a few computers have been specifically designed for the simulation of spin glasses, such as the Ogielski machine [229], SUE [88], and the Janus I [45, 24] and II supercomputers [28].
@@
The great generality of the spin-glass problem uncovered through the computer-science perspective (Sec. I.1) has encouraged people to design and build computers for finding ground states of spin glasses of various types. This includes D-wave’s quantum annealing chips [160, 210], as well as hardware based on a variety of algorithms and/or physical principles [see, e.g., 133, 209, 267, 210, 211].
@@
III.1 Some important numerical results in $D=3$
An overview of analytical and numerical results for the equilibrium spin-glass phase is beyond the scope of this review [for a recent account see, e.g., 208, and references therein]. This section discusses a couple of results that will be of use for the interpretation of non-equilibrium data.
@@ III.1 Some important numerical results in $D=3$
Perhaps the most straightforward quantitative description of the spin-glass phase comes from the probability density function $P(q)$ for the spin overlap $q$ exhibited in Eq. (5). This is plotted in Fig. 19 for several system sizes and a temperature of $T=0.703\approx 0.64T_{\mathrm{g}}$. Notice the two symmetrical peaks at $q=\pm q_{\text{\tiny EA}}(L)$ and, crucially, that the probability density does not go to zero at $q=0$. Many other studies conducted over the years [47, 200, 205, 169, 46, 288] are in agreement on this point. Should $P(q)$ conserve this shape in the large-$L$ limit, the resulting picture would be consistent with RSB —recall Sec. II.2— but this extrapolation is controversial to this day [214, 297, 59, 60, 57, 287]. The dispute is, however, immaterial for our purposes because, in order to model the behavior of non-equilibrium spin-glass dynamics at finite times, we shall only need the value of $P(q)$ for finite sizes.
@@
The authors of Fernandez et al. [111] have studied the chaotic parameter $X^{J}_{T_{1},T_{2}}$, defined by Eq. (13), with $T_{1}=0.703\approx 0.64\,T_{\mathrm{g}}$ and $T_{2}$ in a range extending from $T_{1}$ to above $T_{\mathrm{g}}$. A summary of their results is exhibited in Fig. 21. The upper panel reports the sample average of the chaotic parameter, Eq. (13). It decreases with system size, but is far from being zero (the predicted value for a model expected to have complete temperature chaos).
The central panel in Fig. 21 provides deeper insight. The chaotic parameter is shown for ten samples that have different autocorrelation times (a proxy for the difficulty in thermalizing that particular sample). In some samples, the chaotic parameter changes very smoothly, while in others sharp changes are observed over a small temperature range. Such large sample-to-sample fluctuations are quantified in the bottom panel, where they are shown to increase with system size in the spin-glass phase.
@@
The critical behavior of the EA model is well described by an $n$-component $\phi^{3}$ field theory with a tensorial coupling in the $n\to 0$ limit [143]. In six dimensions (the upper critical dimension in zero magnetic field, recall Sec. II.1), the phase transition exhibits mean-field exponents with computable logarithmic corrections [256, 257]. Near and below $D=6$, there should be a phase transition with non-trivial critical exponents [143, 136]. This scenario has been tested in numerical simulations in dimensions four and six with very good agreement [23, 285, 243, 207, 3]. The low-temperature phase in four dimensions follows the RSB predictions. Both the overlap and the link overlap probability distributions are non-trivial.
The behavior of the EA model in the presence of a magnetic field is not as well understood. A first attempt is to try the Wilsonian approach [292, 14]. One begins with the mean-field effective Hamiltonian [73] and introduces fluctuations to determine at which dimension these fluctuations are relevant. This determines the upper critical dimension in a field, $D^{H}_{\text{u}}$.
This upper critical dimension can also be defined as the smallest dimension in which the critical exponents take the mean-field values. However, in contrast with the zero-field case, where the effective Hamiltonian provides $D_{\text{u}}=6$, no stable fixed point at zero coupling has been found in six dimensions for one-loop computations [73, 248]. The two-loop computation finds that the origin (Gaussian fixed point) is unstable, but finds another (non-trivial and maybe nonperturbative) fixed point that could describe the transition in a field  [84, 82].
@@
The droplet model finds that there is no stable spin-glass phase in a magnetic field at any finite dimension [212, 113, 71, 115]. A modified version of the droplet [293] finds a phase transition only for $D>6$ [see also 72, 247].
@@
As in zero field, numerical simulations in finite dimensions can be supplemented by numerical simulations in one-dimensional lattices with couplings decaying in an algebraic way (the power law of the decay can be related to the dimensions of regular lattices). The results are contradictory: some studies claim that no spin-glass phase in a field survives below six dimensions [170, 168, 278, 279], while other groups find a phase transition [189, 99].
To summarize the situation, it is widely accepted that the upper critical dimension in the absence of a magnetic field is 6. The behavior below six dimensions agrees with the predictions of the RSB theory, at least for the system sizes that one can equilibrate. Instead, there are still many important open issues regarding the equilibrium behavior in an externally applied magnetic field.
@@ IV.1 Aging in experiments
This interpretation was accepted until the experiment by Chamberlin et al. [81], who measured the FC magnetization, then cut the field to zero to extract the time dependence of the TRM. They found that the TRM time decay depended on the time spent in the FC state. In other words, though the FC magnetization was itself time independent, its subsequent dynamics after a magnetic field change depended on the time spent in the FC state.
This demonstrated that, though the FC magnetization may have been time independent, it was anything but an equilibrium state. Lundgren et al. [194], using ZFC time-dependent measurements, showed similar dynamics. The two are complimentary according to the superposition principle first enunciated by the Uppsala group [228, 193, 100] [we reproduce Eq. (1) for the reader’s convenience]:
$M_{\text{TRM}}(t,t_{\mathrm{w}})+M_{\text{FC}}(t,t_{\mathrm{w}})=M_{\text{FC}}(0,t_{\mathrm{w}}+t).$ (20)
We now know that this relationship is valid only in the $H\rightarrow 0$ limit [231, see Sec. V.3]. That analysis uses the Hamming distance, Hd, the “distance” between spin-glass states associated with free-energy barriers $\Delta(t,t_{\mathrm{w}};H)$. It is defined in terms of the overlap between two states, $\alpha$ and $\beta$, with $q_{\alpha,\beta}$ as in Eq. (9), where the state has evolved from its initial quenched state $\alpha(t=0,t_{\mathrm{w}}=0;H)$ to $\beta(t,t_{\mathrm{w}};H)$ in a magnetic field $H$:
@@
The free-energy barrier separating states is thus proportional to their distance Hd, as exhibited in Fig. 22. The specific relationship between Hd and barrier heights was first developed by Vertechi and Virasoro [280].
@@
The significance of Fig. 22 is that, in the presence of a magnetic field $H$, the barriers are quenched for values of Hd such that $\Delta(\text{Hd})\leq|E_{\text{Z}}|\equiv\Delta_{E_{\text{Z}}}(\text{Hd})$ with $E_{\text{Z}}$ defined by Eq. (10). Equation (21) can be written alternatively to take into account this cancellation. Take $q_{\alpha\beta}(E_{\text{Z}})=q_{\text{EA}}-q(E_{\text{Z}})$ to be the value of $q_{\alpha\beta}$ at the value of $q$ where $\Delta(q)=|E_{\text{Z}}|$. The probability of finding a value of $q$ larger than $q(E_{\text{Z}})$ is zero. That is, the states between $q_{\text{EA}}$ and $q(E_{\text{Z}})$ have been emptied after the magnetic field $H$ has been cut to zero in a TRM protocol.
[…]
Lederman et al. [188] found explicit evidence for an upward curvature of $\Delta$ as a function of Hd in a AgMn 2.56 at.spin glass. The calculation of Vertechi and Virasoro [280] finds a similar behavior. It is for this reason that Paga et al. [231] concluded that the principle of superposition is only valid in the $H\rightarrow 0$ limit.
IV.2 Correlation functions for non-equilibrium dynamics and the coherence length

### v2

@@ II.8 Spin-glass simulations and dedicated computers
Interest from computer scientists arises because finding the ground state of a spin glass is an archetypically hard computational problem. Indeed, if the interactions of the spin glass can only be drawn in a nonplanar graph, [^3] A planar graph is one that can be drawn on a piece of paper without edge crossings, such as the square lattice with open boundary conditions. The difference between planar and two-dimensional is important, as discussed below. then finding the ground state belongs to the category of NP-complete problems [42, 184]. Should a truly efficient algorithm for finding the ground state of a nonplanar spin glass be discovered, efficient algorithms for a large category of hard computational problems (such as the traveling salesman problem) would follow [293].
@@
Physicists are often more interested in the behavior of a system at finite temperature $T$ (one could say that physicists wish to minimize the free energy, rather than the energy). However, the physicist and the computer-scientist problems are tightly connected, because the free energy becomes the energy when we approach T=0T\!=\!0.
@@
Difficulties faced by physicists depend strongly on whether they consider equilibrium or off-equilibrium dynamics:
@@
  • Equilibrium studies are (unfortunately) limited to small systems because reaching thermal equilibrium may require an inordinately long simulation time for some samples [in cubic lattices, equilibration times at low temperatures fluctuate from sample to sample by orders of magnitude, 14, 369, 126]. Further computational difficulties are posed by the fact that a large number of samples need to be simulated in order to get meaningful results, because quantities of interest wildly fluctuate from sample to sample.
@@
  • Nonequilibrium studies are conducted on lattices of large linear size $L$, sufficient to ensure that the coherence length $\xi(t,t_{\mathrm{w}};H)$ is always much smaller than $L$. This condition is required in order to compare with laboratory correlation lengths, which, in general, are smaller than physical sample sizes (with the exception of thin films, see Sec. V.1). Indeed, in space dimension $D$, the system can be approximated as composed of $L/\xi(t)]^{D}$ more or less independent samples of effective size $L_{\text{eff}}\approx\xi(t)$. Hence, in close analogy to experiments, a simulation of just one very large sample automatically carries out the average over samples. The main difficulty is that one is limited to simulation algorithms that mimic natural dynamics, such as Metropolis [see, e.g., 328] and, as already explained in Sec. II.6, the natural growth of $\xi(t,t_{\mathrm{w}};H)$ is slow. This slowness calls for very long simulation times on few very large lattices.
@@
Simulating spin glasses is not always hard. Efficient algorithms for finding the ground state are known for planar interaction graphs. It is also important to note that planar and two-dimensional are not quite the same thing: the bilayer that was shown in Barahona [42] to pose an NP-hard problem is certainly two-dimensional but nonplanar. The square lattice with periodic boundary conditions is mildly nonplanar, but still allows one to find ground states for very large lattices [210, 84]. For $T>0$, the problem is also simpler on planar (or quasi-planar) graphs, such as the square lattice with periodic boundary conditions, for which nonlocal simulation algorithms (named cluster algorithms) can shorten simulation times by orders of magnitude [177, 358].
@@
Unfortunately, the situation is much worse for nonplanar graphs, such as the simple-cubic lattice used in most $D=3$ simulations. At T=0T\!=\!0, ground states have been found in cubic lattices with size $L\!\leq\!14$ with the help of a heuristic algorithm [238]. See Caracciolo et al. [84] for information on sure solvers. Many algorithms have been proposed for T>0T\!>\!0: simulated tempering [237], umbrella sampling [see, e.g., 50, 185], the Wang-Landau algorithm [355] or population annealing [see, e.g., 235]. Nevertheless, the best-performing algorithm for T>0T\!>\!0 equilibrium simulations appears to be parallel tempering [152, 178, 339], which, with the help of a custom-built computer, has been able to equilibrate cubic lattices of size L=32L\!=\!32 at $T\!\approx\!0.64T_{\mathrm{g}}$ [15]. Temperature chaos, Sec. II.4, is the physical mechanism that makes it difficult for algorithms that exploit the correlation at different temperatures (e.g., simulated annealing or parallel tempering) to reach thermal equilibrium at even lower temperatures [128, 61]. [^4] In a very chaotic sample, thermalization achieved at a higher temperature would be almost irrelevant for sampling at a lower temperature. Thus, the thermalization process would need to restart at the lower temperature.
@@
Lacking better algorithms, the best we can do is to increase computing power. Computers specifically designed for high-energy-physics simulations, such as RTN [24] and APE100 [44], have also been quite useful for spin glasses [see Badoni et al. [25], Ciria et al. [98] in the case of RTN and Marinari et al. [245, 244, 240] for APE100]. More recently, graphics processing units (GPUs) have proven to be extremely powerful for simulating quantum spin glasses [52]. Furthermore, a few computers have been specifically designed for the simulation of spin glasses, such as the Ogielski machine [285], SUE [101], and the Janus I [47, 26] and II supercomputers [30].
@@
The great generality of the spin-glass problem uncovered through the computer-science perspective (Sec. I.1) has encouraged people to design and build computers for finding ground states of spin glasses of various types. This includes D-wave’s quantum annealing chips [189, 252], as well as hardware based on a variety of algorithms and/or physical principles [see, e.g., 157, 250, 334, 252, 253].
@@
[…]
An overview of analytical and numerical results for the equilibrium spin-glass phase is beyond the scope of this review [for a recent account see, e.g., 248, and references therein]. This section discusses some results that will be of use for the interpretation of nonequilibrium data.
@@
Perhaps the most straightforward quantitative description of the spin-glass phase comes from the probability density function $P(q)$ for the spin overlap $q$ exhibited in Eq. (5). This is plotted in Fig. 20 for several system sizes and a temperature of $T=0.703\approx 0.64T_{\mathrm{g}}$. Notice the two symmetrical peaks at $q=\pm q_{\text{\tiny EA}}(L)$ and, crucially, that the probability density does not go to zero at $q=0$. Many other studies conducted over the years [49, 239, 244, 198, 48, 360] are in agreement on this point. Should $P(q)$ conserve this shape in the large-$L$ limit, the resulting picture would be consistent with RSB —recall Sec. II.2— but this extrapolation is controversial to this day [262, 372, 62, 63, 60, 359]. The dispute is, however, immaterial for our purposes because, in order to model the behavior of nonequilibrium spin-glass dynamics at finite times, we shall only need the value of $P(q)$ for finite sizes.
@@
The authors of Fernandez et al. [128] have studied the chaoticity parameter $X^{J}_{T_{1},T_{2}}$, defined by Eq. (10), with $T_{1}=0.703\approx 0.64\,T_{\mathrm{g}}$ and $T_{2}$ in a range extending from $T_{1}$ to above $T_{\mathrm{g}}$. A summary of their results is exhibited in Fig. 22. The upper panel reports the sample average of the chaoticity parameter, Eq. (10). It decreases with system size, but is far from being zero (the predicted value for a model expected to have complete temperature chaos).
The central panel in Fig. 22 provides deeper insight. The chaoticity parameter is shown for ten samples that have different autocorrelation times (a proxy for the difficulty in thermalizing that particular sample). In some samples, the chaoticity parameter changes very smoothly, while in others sharp changes are observed over a small temperature range. Such large sample-to-sample fluctuations are quantified in the bottom panel, where they are shown to increase with system size in the spin-glass phase.
@@
The critical behavior of the EA model is well described by an $n$-component $\phi^{3}$ field theory with a tensorial coupling in the $n\to 0$ limit [167]. In six dimensions (the upper critical dimension in zero magnetic field, recall Sec. II.1), the phase transition exhibits mean-field exponents with computable logarithmic corrections [321, 322]. Near and below $D=6$, there should be a phase transition with nontrivial critical exponents [167, 160]. This scenario has been tested in numerical simulations in dimensions four and six with very good agreement [25, 357, 301, 246, 3]. The low-temperature phase in four dimensions follows the RSB predictions. Both the overlap and the link overlap probability distributions are nontrivial.
The behavior of the EA model in the presence of a magnetic field is not well understood. A first attempt is to try the Wilsonian approach [367, 16]. One begins with the mean-field effective Hamiltonian [78] and introduces fluctuations to determine at which dimension these fluctuations are relevant. This determines the upper critical dimension in a field, $D^{H}_{\text{u}}$.
This upper critical dimension can also be defined as the smallest dimension in which the critical exponents take the mean-field values. However, in contrast with the zero-field case, where the effective Hamiltonian provides $D_{\text{u}}=6$, no stable fixed point at zero coupling has been found in six dimensions for one-loop computations [78, 311]. The two-loop computation finds that the origin (Gaussian fixed point) is unstable, but finds another (nontrivial and maybe nonperturbative) fixed point that could describe the transition in a field  [96, 94].
@@
The droplet model finds that there is no stable spin-glass phase in a magnetic field at any finite dimension [254, 134, 76, 136]. A modified version of the droplet [368] finds a phase transition only for $D>6$ [see also 77, 305].
@@
As in zero field, numerical simulations in finite dimensions can be supplemented by numerical simulations in one-dimensional lattices with couplings decaying algebraically (the power law of the decay can be related to the dimensions of regular lattices). The results are contradictory: some studies claim that no spin-glass phase in a field survives below six dimensions [199, 197, 348, 349], while other groups find a phase transition [226, 114].
To summarize the situation, it is widely accepted that the upper critical dimension in the absence of a magnetic field is 6. The behavior below six dimensions agrees with the predictions of the RSB theory, at least for system sizes that one can equilibrate. However, there remain many important open issues regarding the equilibrium behavior in an externally applied magnetic field.
@@ IV.1 Aging in experiments
The concept of aging is not exclusive to spin glasses. It is a ubiquitous phenomenon found in a wide variety of materials [204, 308]. Examples are polymer mechanics [333]; structural glasses [331, 315]; strongly correlated electronic systems [356]; gelatin gels [306]; magnetic nanoparticle assemblies [307]; and hysterons, models of interacting binary systems, from crumpled sheets to frustrated metamaterials [171].
@@
This interpretation was accepted until the experiment by Chamberlin et al. [93], who measured the FC magnetization, then cut the field to zero to extract the time dependence of the TRM. The TRM time decay was found to depend on the time spent in the FC state. In other words, though the FC magnetization was itself approximately time independent, its subsequent dynamics after a magnetic field change depended on the time spent in the FC state.
This demonstrated that the FC magnetization was anything but an equilibrium state. Lundgren et al. [233], using ZFC time-dependent measurements, showed similar dynamics. The two are complimentary according to the superposition principle first enunciated by the Uppsala group [283, 232, 115] [we reproduce Eq. (1) for the reader’s convenience]:
$M_{\text{TRM}}(t,t_{\mathrm{w}})+M_{\text{ZFC}}(t,t_{\mathrm{w}})=M_{\text{FC}}(0,t_{\mathrm{w}}+t).$ (22)
We now know that this relationship is valid only in the $H\rightarrow 0$ limit [288, see Sec. V.3]. That analysis uses the Hamming distance, Hd, the “distance” between spin-glass states associated with free-energy barriers $\Delta(t,t_{\mathrm{w}};H)$. It is defined in terms of the overlap between two states, $\alpha$ and $\beta$, with $q_{\alpha,\beta}$ as in Eq. (9), where the state has evolved from its initial quenched state $\alpha(t=0,t_{\mathrm{w}}=0;H)$ to $\beta(t,t_{\mathrm{w}};H)$ in a magnetic field $H$:
@@
The free-energy barrier separating states has been found to increase more rapidly than linear with Hd [288] as exhibited in Fig. 23. The specific relationship between Hd and barrier heights was first developed by Vertechi and Virasoro [350].
@@
The significance of Fig. 23 is that, in the presence of a magnetic field $H$, the barriers are quenched for values of Hd such that $\Delta(\text{Hd})\leq|E_{\text{Z}}|\equiv\Delta_{E_{\text{Z}}}(\text{Hd})$ with $E_{\text{Z}}$ defined by Eq. (13). Equation (23) can be written alternatively to take into account this cancellation. Take $q_{\alpha\beta}(E_{\text{Z}})=q_{\text{EA}}-q(E_{\text{Z}})$ to be the value of $q_{\alpha\beta}$ at the value of $q$ where $\Delta(q)=|E_{\text{Z}}|$. The probability of finding a value of $q$ larger than $q(E_{\text{Z}})$ is zero. That is, the states between $q_{\text{EA}}$ and $q(E_{\text{Z}})$ have been immediately emptied after the magnetic field $H$ was cut to zero in a TRM protocol.
The spin-glass coherence length $\xi(t=0,t_{\mathrm{w}};H)$ therefore grows not from $q=0$ but from $q=q(E_{\text{Z}})$ in a TRM experiment. If $\Delta(t_{\mathrm{w}})$ would depend linearly on Hd, equivalently on $q_{\alpha\beta}$ from Eq. (23), there would be no difference in $\xi(t=0,t_{\mathrm{w}};H)$ between a TRM and a ZFC experiment. However, if instead $\Delta(t_{\mathrm{w}})$ behaves as drawn in Fig. 23, that is, exhibits an upward curvature as Hd increases, a TRM protocol would begin at a larger value of $q$ than the ZFC protocol which begins at $q=0$. The more rapid increase of $\Delta(t_{\mathrm{w}})$ as Hd increases from Hd($\Delta=|E_{Z}|)$, as compared to increasing from Hd($\Delta=0)$, would imply a slower growth of $\xi(t,t_{\mathrm{w}};H)$ for a TRM protocol than for an ZFC protocol. Hence, a breakdown of the superposition principle.
Lederman et al. [224] found explicit evidence for an upward curvature of $\Delta$ as a function of Hd in a AgMn 2.56 at.% spin glass. The calculation of Vertechi and Virasoro [350] found a similar behavior. It is for this reason that Paga et al. [288] concluded that the principle of superposition only holds in the $H\rightarrow 0$ limit.
IV.2 Correlation functions for nonequilibrium dynamics and the coherence length